OpenStudy (anonymous):

Find the direct variation equation of the graph through the points (0, 0) and (1, -2). Write in y=kx form. Answer y = 2x y = -2x y = 1/2 x y = -1/2 x

OpenStudy (anonymous):

here is an example of direct variation! Example A: If y varies directly as x, and y = 8 when x = 12, find k and write an equation that expresses this variation. Plan of Attack: Plug the given values into the equation y = kx. Solve for k. Then replace k with its value in the equation y = kx. Step-by-Step: Start with our standard equation: y = kx Insert our known values: 8 = k*12 Divide both sides by 12 to find k: 8/12 = k 2/3 = k Next: Go back to y = kx and replace k with 2/3. y = (2/3)x

OpenStudy (anonymous):

so just take this and plug in the points that ju have

OpenStudy (anonymous):

so it would be y=-2x?

OpenStudy (anonymous):

start from the beginning lol plug in the 1 for y and the -2 for x then start to solve for k

OpenStudy (anonymous):

1=k(-2) then divide by -2 on both sides

OpenStudy (anonymous):

1=k*(-2) divid it by -2 an k = -1/2

OpenStudy (anonymous):

yesh then plug k in and there is ju answer! y=(-1/2)x

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

welcome! dang it seems like i know everything... but i dont lmao